Proposal Year 0 Year 3 IRR Year 1 Year 2 60.0% $30 $153 $88 A -$100 $206 $95 55.6% $37 50.2% -$100 S0

Find the missing values and calculate NPV for proposals A, B, and C. What proposal should be chosen?
Answer #1

Proposal A has all the values

Proposal B:

IRR is the rate of return that makes Cash outflow equal to present value of cash inflow

Year 0 = 0 + 206 / (1 + 0.556)2 + 95 / (1 + 0.556)3

Year 0 = 0 + 85.08 + 25.217

Year 0 = -$110

Proposal C:

100 = 37 / (1 + 0.502)1 + 0 + Year 3 / (1 + 0.502)3

100 = 24.634 + Year 3 / 3.38852

Year 3 = 255

Discount rate is missing in the question. If you have the discount rate, please put it in comments

NPV:

Proposal A:

NPV = Present value of cash inflows - present value of cash outflows

NPV = -100 + 30 / (1 + 0.1)1 + 153 / (1 + 0.1)2 + 88 / (1 + 0.1)3

NPV = -100 + 27.27 + 126.45 + 66.12

NPV = $120

Proposal B:

NPV = Present value of cash inflows - present value of cash outflows

NPV = -110 + 0 / (1 + 0.1)1 + 206 / (1 + 0.1)2 + 95 / (1 + 0.1)3

NPV = -110 + 0 + 170.25 + 71.37

NPV = $132

Proposal C:

NPV = Present value of cash inflows - present value of cash outflows

NPV = -100 + 37 / (1 + 0.1)1 + 0 / (1 + 0.1)2 + 255 / (1 + 0.1)3

NPV = -100 + 33.64 + 0 + 191.59

NPV = $125

Proposal B should be chosen as it has the highest NPV